Consider the following information for a particular economy.
Total population =60 million
Number of employed=30 million
Total labor force = 40 million
Natural rate of unemployment =12%
a) Find the total unemployment rate
b) Calculate the cyclical unemployment rate
1
Expert's answer
2020-05-28T11:01:56-0400

Unemployment rate = Number of Unemployed Persons / Labor Force

Unemployment rate = (40-30)/40 = 25%

Cyclical unemployment rate = Total unemployment rate - Natural rate of unemployment

Cyclical unemployment rate = 25-12= 13%